OpenStudy (anonymous):

write the slope-intercept equation for the line with slope 1/2 and y-intercept (0,-5)

OpenStudy (lgbasallote):

use the equation \[\huge y - y_1 = m (x -x_1)\] m is the slope

OpenStudy (anonymous):

y=1/2x-5

OpenStudy (anonymous):

would it be y=1/2x- -5????

OpenStudy (anonymous):

are you asking if it's --5?

OpenStudy (anonymous):

yes im asking if it is negative 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you